Given the following data, what is the average cycle time? Assume for each story, it starts at 8:00 AM on the start day and ends at 8:00 AM on the end day.

Story

Story Start Day

Story End Day

Story Duration

1

Day 1

Day 2

2

Day 3

Day 4

3

Day 4

Day 8

4

Day 6

Day 9

5

Day 5

Day 10
Answer #1

Story 1: Starts at 8:00 AM Day 1 and Ends at 8:00 AM Day 2 = Story duration is 1 Day

Story 2: Starts at 8:00 AM Day 3 and Ends at 8:00 AM Day 4 = Story duration is 1 Day

Story 3: Starts at 8:00 AM Day 4 and Ends at 8:00 AM Day 8 = Story duration is 4 Days

Story 4: Starts at 8:00 AM Day 6 and Ends at 8:00 AM Day 9 = Story duration is 3 Days

Story 5: Starts at 8:00 AM Day 5 and Ends at 8:00 AM Day 10 = Story duration is 5 Days

Total days = 10 days

Total story duration = (1 + 1 + 4 + 3 + 5) = 14 days

5 stories are completed in 10 days.(if the stories are not started continuously).

So the average cycle time 10/5 days per story = 2 days per story or 5/10 stories per day = 0.5 stories per day.

If the stories are started one after another, then 5 stories are completed in 14 days.

The average cycle when the stories are started continuously one after another then, average cycle time = 14/5 days per story = 2.8 days per story or 5/14 stories per day = 0.36 stories per day